Find the x-intercepts:
y = x^2 - 3x - 1
How do I do this problem? I am confused on this one completely! Thank you for your help!

Answer by stanbon(75887)

Find the x-intercepts:
y = x^2 - 3x - 1
---------------
To find the x-intercepts (roots), let y=0
x^2-3x-1=0
Use the quadratic formula to find "x":
x=[3+-sqrt(9-4*-1]/2
x=(3+sqrt(13))/2 or x=(3-sqrt(13))/2
x=0.3027756... or x=3.3027756...
---------
Cheers,
Stan H.